An object starts from rest with an acceleration of 5m/s2 along a straight line. Find the speed and distance travelled after 6 seconds what use the average speed during those
6 seconds?
1
Expert's answer
2020-09-05T14:44:29-0400

Let initial velocity (v0)=0 m/s

Acceleration (a)=5 m/s2

Velocity after 6 seconds

v6=vo+a x t =5x6=30 m/s

Velocity after 5 seconds

v5=vo+a x t =5x5=25 m/s

Average speed during 6th second

=1/2(v5+v6)=27.5 m/s

Distance travelled after 6 second

d=vo x t +1/2(a)x t2 =90 m